User Tools

Site Tools


lbaops:lbajul2011:v252yholog

Agilent 16.6 GHz level 16 dBm

SML LO to translator 636.0 MHz ext ref 10.0 dBm

VSOP_HO.pro das profile

Recording channels 0,1,2,3 to ATNF V016B raid.

Statistics look good. But I cannot find the coherence test tone. Fringes were okay on the fringe test source vc145.

Temp 6.9C, pressure 1027.6 HPa, humidity 82.3%, wind 10.9 km/h at UT 16:40

Frequency counter reading 0.01702

fmout-gps = +20.4 microseconds

John Dickey observer

At around 2330, Noise diode problem was fixed. The exact time will be in the FS log. The noise diode control had not been set to remote before this and so Tsys measurements in the FS log would not have been working for the first few hours of the experiment.

FS time was corrected after being 1 second fast. This problem would not have affected pointing but only “source=” commands.

Brett Reid.

Stopped recording about 5 min early, no other problems. (fringes found).

Claire Hotan

lbaops/lbajul2011/v252yholog.txt · Last modified: 2015/12/18 16:38 (external edit)